# short-time Fourier transform iSTFT inverse short-time Fourier transform

### v1

@@ V-A Speech Dereverberation
TABLE II: Speech dereverberation results on reverberant VCTK datasets. We indicate for each method in the table whether it is supervised or not. Boldface numbers indicate best performance for supervised and unsupervised methods separately.
Matched Mismatched
Method Unsup. DNS-MOS PESQ ESTOI DNS-MOS PESQ ESTOI
Reverberant - 3.14 $\pm$ 0.52 1.61 $\pm$ 0.37 0.50 $\pm$ 0.14 3.05 $\pm$ 0.47 1.57 $\pm$ 0.29 0.47 $\pm$ 0.11
PSE 3.75 $\pm$ 0.38 2.85 $\pm$ 0.55 0.80 $\pm$ 0.10 3.61 $\pm$ 0.39 2.08 $\pm$ 0.47 0.64 $\pm$ 0.09
SGMSE+M [24, 55] 3.88 $\pm$ 0.32 2.99 $\pm$ 0.48 0.78 $\pm$ 0.09 3.74 $\pm$ 0.34 2.48 $\pm$ 0.47 0.69 $\pm$ 0.09
StoRM [25] 3.90 $\pm$ 0.33 3.33 $\pm$ 0.48 0.82 $\pm$ 0.10 3.83 $\pm$ 0.32 2.51 $\pm$ 0.53 0.67 $\pm$ 0.09
Yohena and Yatabe [8] 2.99 $\pm$ 0.56 1.80 $\pm$ 0.33 0.55 $\pm$ 0.12 2.94 $\pm$ 0.44 1.71 $\pm$ 0.29 0.51 $\pm$ 0.10
WPE [56] 3.24 $\pm$ 0.54 1.81 $\pm$ 0.42 0.57 $\pm$ 0.14 3.10 $\pm$ 0.48 1.74 $\pm$ 0.37 0.54 $\pm$ 0.12
Saito et al. [34] 3.22 $\pm$ 0.56 1.68 $\pm$ 0.40 0.51 $\pm$ 0.13 3.12 $\pm$ 0.52 1.70 $\pm$ 0.33 0.52 $\pm$ 0.10
GibbsDDRM [35] 3.33 $\pm$ 0.53 1.70 $\pm$ 0.37 0.51 $\pm$ 0.13 3.30 $\pm$ 0.52 1.75 $\pm$ 0.36 0.52 $\pm$ 0.11
RVAE-EM [38] 3.05 $\pm$ 0.53 1.83 $\pm$ 0.32 0.54 $\pm$ 0.11 3.00 $\pm$ 0.45 1.76 $\pm$ 0.30 0.52 $\pm$ 0.10
BUDDy (ours) 3.76 $\pm$ 0.41 2.30 $\pm$ 0.53 0.66 $\pm$ 0.12 3.74 $\pm$ 0.38 2.24 $\pm$ 0.54 0.65 $\pm$ 0.12
V-A1 Data
We use VCTK [57] as clean speech, selecting 103 speakers for training, 2 for validation and 2 for testing. The total dataset represents 44h of audio, which we down-sample to 16kHz for our experiments. We curate RIRs from various public datasets [58, 59, 60, 61, 62, 63, 64, 65, 66]. In total we approximately obtain 10k RIRs, and split them between training, validation and testing using ratios 0.9/0.05/0.05. We also generate another RIR dataset for testing methods in a mismatched setting, using RIRs simulated with pyroomacoustics[67]. For ease of comparison, we choose simulation parameters such that the distributions of reverberation times and direct-to-reverberation ratios of the simulated mismatched dataset approximately match those of the real-recorded matched dataset.
V-A2 Baselines
@@ V-A4 Instrumental metrics
For instrumental evaluation of the speech dereverberation performance, we use the intrusive Perceptual Evaluation of Speech Quality (PESQ) [69] and extended short-term objective intelligibility (ESTOI) [70] for assessment of speech quality and intelligibility respectively. We also use the non-intrusive DNS-MOS [71], a DNN-based mean opinion score (MOS) approximation following the ITU-T P.835 recommendation [72].
V-A5 Instrumental evaluation results
@@
We display in Table II the dereverberation results for all blind methods, both supervised and unsupervised. Blind supervised approachesPSE, SGMSE+ and StoRM generally perform better than unsupervised methods as they benefit from supervision at training time. However, we can observe the limited generalization ability of supervised approaches in the setting where acoustic conditions are mismatched, i.e. when using simulated RIRs. Our method BUDDy, however, seamlessly adapts to changing acoustics since it was trained without supervision. This enables BUDDy to keep the same performance on both matched and mismatched datasets, where supervised methods like PSE lose as much as 0.77 PESQ points in mismatched conditions. Furthermore, BUDDy performs far better than all other blind unsupervised baselines. For instance, BUDDy outperforms WPE by as much as 0.50 PESQ and 0.10 ESTOI points. Indeed, traditional unsupervised methods [7, 8] only draw limited benefits from their uninformed Gaussian prior on anechoic speech, while diffusion-based Saito et al. [34] and GibbsDDRM [35] seem to only marginally deviate from their WPE initialization. RVAE-EM [38] also obtains low instrumental scores, but informal listening suggested that its dereverberation abilities were superior to those of WPE.
V-A6 Ablation study
@@
We conduct an ablation study to evaluate the impact of the projection step (13) introduced in the operator optimization (see Section IV-A). We present the results in Table III and observe that, although the minimum-phase consistency projection has a theoretical justification as a mean to enhance the stability of the inverse RIR during optimization, its practical effect appears negligible. However, we observe that the other operations in the projection step, i.e. STFT consistency, enforcement of the direct path, and speech magnitude constraint, are all instrumental in guiding BUDDy toward a solution with higher fidelity to clean speech, as measured by PESQ. We show DNS-MOS figures out of completeness. However, DNS-MOS variations are small across ablations and not indicative of fidelity to reference speech as DNS-MOS is not an intrusive metric.
@@
Additionally, we examine the effect of parameterizing the likelihood model with a $L^{2}$-distance on compressed spectrograms rather than on waveforms as in previous work [26]. To do so, we replace the cost function $C(\cdot,\cdot)$ from (8), which is based on compressed spectrogams, with a simpler waveform-domain $L^{2}$-distance. The results clearly show the superiority of using a cost function on compressed spectrograms.
V-A7 Listening experiment
@@
Instrumental metrics offer only limited insights into the performance of dereverberation algorithms [73]. We therefore conduct a listening experiment based on the MUSHRA recommendation [74] to assess the performance of BUDDy as perceived by human listeners. The test comprised 12 pages, featuring 6 reverberant speech utterances from each of the matched and mismatched datasets. Participants were asked to rate the different stimuli with a single number representing overall quality, taking into account factors such as voice distortion, residual reverberation, and potential artifacts [73]. The test stimuli include our proposed method BUDDy, the unsupervised WPE [7] and RVAE-EM [38], as well as the supervised baselines PSE and SGMSE [24], Further details on the organization of the listening experiment are reported in Appendix -C1.
TABLE III: Ablation study on reverberant VCTK dataset.
@@
- RMS Power Constraint 2.22 $\pm$ 0.50 3.64 $\pm$ 0.50
- Fixed Direct Path 2.10 $\pm$ 0.46 3.78 $\pm$ 0.44
- STFT Consistency 1.96 $\pm$ 0.41 3.84 $\pm$ 0.39
@@
The results of the experiment are presented in Fig. 3. It can be observed that the unsupervised baselines WPE and RVAE-EM received low scores. Yet, RVAE-EM performs consistently better than WPE in this listening experiment, as opposed to what is suggested by instrumental metrics in Table II. In the matched test set (Fig. 3a), BUDDy obtained significantly lower scores than PSE and SGMSE+ ($p<0.001$ in a paired Welch test). However, in the mismatched set, PSE and SGMSE+ suffered a decrease in performance, losing up to $20$ points (out of $100$), while BUDDy mantained similar scores. In that case, there is no significant difference in performance between the three approaches ($p>0.1$), which closes the gap between BUDDy and the top-performing supervised baselines in this mismatched setting, highlighting the advantage provided by unsupervised learning.
Figure 3: Listening test results on reverberant VCTK datasets. The boxplot shows first quartile, median and third quartile.
V-B Singing Voice Dereverberation
@@
TABLE IV: Singing voice dereverberation results on reverberant NHSS dataset. We indicate each method whether it is unsupervised or not.
Method Unsup. $\ell^{1}$ STFT FAD (VGGish)
Reverberant - 1.98 $\pm$ 0.66 6.41
Conditional Diffusion 1.56 $\pm$ 0.48 1.71
WPE [7] 2.02 $\pm$ 0.65 4.53
Saito et al. [34] 1.95 $\pm$ 0.65 5.41
BUDDy (ours) 1.92 $\pm$ 0.60 1.32
@@
Figure 4: Listening test results on singing voice dataset. The boxplot shows first quartile, median and third quartile.
V-B1 Data
@@
We collect several publicly available singing voice datasets [75, 76, 77, 78, 79, 80]. These datasets feature over 94 h of studio-quality solo singing from a diverse array of singers and singing styles, spanning various languages. The majority of the recordings are in Chinese, followed by English, Japanese, and Korean. All datasets are down-sampled to 44.1 kHz. For testing, similar to [35], we use the sung part of NHSS [81, 82]. The NHSS dataset contains 100 English-language pop songs, 10 for each of the five male and five female singers recruited. We select a subset (90%) of the RIRs curated for the VCTK-based experiments, such that we only retain the RIRs whose original sample rate is at least 44.1kHz.
V-B2 Baselines
We evaluate the performance of BUDDy against two unsupervised baselines: WPE [7] and the unsupervised method from Saito et al. [34] which was originally designed for singing voice dereverberation. Due to the lack of established supervised baselines for this specific task, we created our own by adapting the diffusion model from BUDDy to a supervised setting. In this adaptation, we trained a diffusion model on dry singing voice using the same architecture and hyperparameters as the unconditional model in BUDDy, but added a paired sample of reverberant voice as a condition. The conditioning sample is incorporated into the architecture by stacking it in the channel dimension. This setup resembles the ”variance exploding” approach used by Gonzalez et al. for speech enhancement [83]. We train all methods on our singing voice dataset detailed in the previous section.
V-B3 Hyperparameters
As in [32] for music restoration, we use the UNet architecture proposed in [84] without self-attention blocks, and wrap the computations within an invertible Constant-Q Transform (CQT) [85]. The CQT produces a time-frequency representation where pitch transpositions are equivalent to frequency-wise translations, highlighting its interest for processing signals with harmonic components such as singing voice and music. More details with regard to the architecture and specific training configuration and inference hyperparameters are reported in Appendix -C2
V-B4 Evaluation metrics

### v2

@@ IV-A Speech Dereverberation
TABLE I: Speech dereverberation results on reverberant VCTK datasets. We indicate for each method in the table whether it is supervised or not. Boldface numbers indicate best performance for supervised and unsupervised methods separately.
VCTK-RealReverb (Matched) VCTK-SimulatedReverb (Mismatched)
Method Unsup. DNS-MOS PESQ ESTOI SI-SDR DNS-MOS PESQ ESTOI SI-SDR
Reverberant - 3.14 $\pm$ 0.52 1.61 $\pm$ 0.37 0.50 $\pm$ 0.14 -12.3 $\pm$ 6.9 3.05 $\pm$ 0.47 1.57 $\pm$ 0.29 0.47 $\pm$ 0.11 -12.5 $\pm$ 8.6
PSE 3.75 $\pm$ 0.38 2.85 $\pm$ 0.55 0.80 $\pm$ 0.10 8.5 $\pm$ 6.5 3.61 $\pm$ 0.39 2.08 $\pm$ 0.47 0.64 $\pm$ 0.09 -8.4 $\pm$ 8.6
SGMSE+M [24, 59] 3.88 $\pm$ 0.32 2.99 $\pm$ 0.48 0.78 $\pm$ 0.09 0.2 $\pm$ 9.3 3.74 $\pm$ 0.34 2.48 $\pm$ 0.47 0.69 $\pm$ 0.09 -8.4 $\pm$ 8.8
StoRM [25] 3.90 $\pm$ 0.33 3.33 $\pm$ 0.48 0.82 $\pm$ 0.10 9.5 $\pm$ 6.5 3.83 $\pm$ 0.32 2.51 $\pm$ 0.53 0.67 $\pm$ 0.09 -8.8 $\pm$ 10.2
Yohena and Yatabe [8] 2.99 $\pm$ 0.56 1.80 $\pm$ 0.33 0.55 $\pm$ 0.12 -11.4 $\pm$ 7.7 2.94 $\pm$ 0.44 1.71 $\pm$ 0.29 0.51 $\pm$ 0.10 -11.4 $\pm$ 8.6
WPE [60] 3.24 $\pm$ 0.54 1.81 $\pm$ 0.42 0.57 $\pm$ 0.14 -11.5 $\pm$ 8.2 3.10 $\pm$ 0.48 1.74 $\pm$ 0.37 0.54 $\pm$ 0.12 -11.4 $\pm$ 8.8
Saito et al. [34] 3.22 $\pm$ 0.56 1.68 $\pm$ 0.40 0.51 $\pm$ 0.13 -11.7 $\pm$ 9.2 3.12 $\pm$ 0.52 1.70 $\pm$ 0.33 0.52 $\pm$ 0.10 -11.7 $\pm$ 8.5
GibbsDDRM [35] 3.33 $\pm$ 0.53 1.70 $\pm$ 0.37 0.51 $\pm$ 0.13 -11.9 $\pm$ 8.5 3.30 $\pm$ 0.52 1.75 $\pm$ 0.36 0.52 $\pm$ 0.11 -11.8 $\pm$ 8.9
RVAE-EM [38] 3.05 $\pm$ 0.53 1.83 $\pm$ 0.32 0.54 $\pm$ 0.11 -12.2 $\pm$ 8.1 3.00 $\pm$ 0.45 1.76 $\pm$ 0.30 0.52 $\pm$ 0.10 -11.8 $\pm$ 8.5
BUDDy (ours) 3.76 $\pm$ 0.41 2.30 $\pm$ 0.53 0.66 $\pm$ 0.12 -7.8 $\pm$ 8.40 3.74 $\pm$ 0.38 2.24 $\pm$ 0.54 0.65 $\pm$ 0.12 -8.4 $\pm$ 9.9
IV-A1 Data
We use VCTK [61] as clean speech, selecting 103 speakers for training, two for validation, and two for testing. The total dataset represents 44 h of audio, which we down-sample to 16 kHz for our experiments. We curate RIRs from various public datasets [62, 63, 64, 65, 66, 67, 68, 69, 70]. In total we approximately obtain 10k RIRs, and split them between training, validation and testing using ratios 0.9/0.05/0.05. We use the three following test benchmarks for our experiments:
[…]
  • VCTK-SimulatedReverb: This test set uses the same test speech utterances as VCTK-RealReverb, matching the training speech corpus. However, the reverberant conditions are obtained by simulating RIRs with pyroomacoustics[71]. For ease of comparison, we choose simulation parameters such that the distributions of reverberation times and direct-to-reverberation ratios of the simulated mismatched dataset approximately match those of the matched dataset using real RIRs.
[…]
IV-A2 Baselines
@@ IV-A4 Instrumental metrics
For instrumental evaluation of the speech dereverberation performance, we use the intrusive Perceptual Evaluation of Speech Quality (PESQ) [75] and extended short-term objective intelligibility (ESTOI) [76] for assessment of speech quality and intelligibility respectively. We also use the non-intrusive DNS-MOS [77], a DNN-based mean opinion score (MOS) approximation following the ITU-T P.835 recommendation [78]. We also report SI-SDR [79], however it must be noted that such point-wise distance metrics do not represent well the performance of generative models, given the natural variability of the corresponding estimates.
IV-A5 Instrumental evaluation results
@@
We display in Table I the dereverberation results for all blind methods, both supervised and unsupervised. Blind supervised approaches PSE, SGMSE+ and StoRM generally perform better than unsupervised methods as they benefit from supervision at training time. However, we can observe the limited generalization ability of supervised approaches on the VCTK-SimulatedReverb when reverberant conditions are not the same as those presented during training. Our method BUDDy, however, seamlessly adapts to changing acoustics since it was trained without supervision. This enables BUDDy to retain its performance from VCTK-RealReverb to VCTK-SimulatedReverb, where supervised methods like PSE lose up to 0.77 PESQ points in mismatched reverberant conditions.
When evaluating on the DDS-DAPS-RealRecorded dataset, as reported in Table II, the generalization gap between our method and the supervised baselines increases even further. On this benchmark, BUDDy outperforms the best supervised baselines SGMSE+ and StoRM on DNS-MOS and ESTOI and has similar PESQ scores. This experiment highlights several points. First, BUDDy is naturally robust to background noise although it is not present in its signal model, echoing the conclusions of our prior work [26]. Informal listening suggests that the background noise is left untouched by the algorithm, and therefore it can be easily dealt with after processing. Furthermore, BUDDy’s unconditional diffusion model trained on the anechoic VCTK dataset seems to generalize well to DAPS speech utterances. Finally, BUDDy also performs strong dereverberation when handling a realistic reverberant model, since the utterances in the DDS-DAPS-RealRecorded dataset are directly recorded in a reverberant room, and not produced by a convolution between an utterance and a measured RIR. On the contrary, the supervised baselines suffer from the multiple mismatches between the presented benchmark and the training conditions, be it with regard to background noise, different speech corpora or reverberant model. This strengthens the position of BUDDy as a robust unsupervised baseline versus top-performing supervised baselines.
Overall, BUDDy performs far better than all other blind unsupervised baselines. For instance, BUDDy outperforms RVAE-EM by as much as 0.47 PESQ and 0.12 ESTOI points. Indeed, traditional unsupervised methods [7, 8] only draw limited benefits from their uninformed Gaussian prior on anechoic speech, while diffusion-based Saito et al. [34] and GibbsDDRM [35] seem to only marginally deviate from their WPE initialization. RVAE-EM [38] also obtains low instrumental scores, but informal listening suggested that its dereverberation abilities were superior to those of WPE.
TABLE II: Speech dereverberation results on DDS-DAPS-RealRecorded reverberant benchmark. Boldface numbers indicate best performance for supervised and unsupervised methods separately. Underlined numbers indicate best performance across all methods.
Method Unsup. DNS-MOS PESQ ESTOI
Reverberant - 2.35 $\pm$ 0.59 1.30 $\pm$ 0.16 0.63 $\pm$ 0.10
PSE 2.91 $\pm$ 0.55 1.65 $\pm$ 0.40 0.74 $\pm$ 0.14
SGMSE+M [24] 3.21 $\pm$ 0.52 2.14 $\pm$ 0.42 0.83 $\pm$ 0.09
StoRM [25] 3.48 $\pm$ 0.45 2.12 $\pm$ 0.50 0.82 $\pm$ 0.12
WPE [60] 2.64 $\pm$ 0.59 1.48 $\pm$ 0.29 0.70 $\pm$ 0.11
RVAE-EM [38] 2.68 $\pm$ 0.55 1.59 $\pm$ 0.33 0.71 $\pm$ 0.11
BUDDy (ours) 3.55$\pm$ 0.49 2.11 $\pm$ 0.47 0.86 $\pm$ 0.10
IV-A6 Ablation study
@@
We conduct an ablation study to evaluate the impact of the projection step (15) introduced in the operator optimization (see Section III-A). We present the results in Table III and observe that, although the minimum-phase consistency projection has a theoretical justification as a mean to enhance the stability of the inverse RIR during optimization, its practical effect appears negligible, which can be due to a mismatch with the fact that real RIRs are generally mixed-phase filters [12]. However, we observe that the other operations in the projection step, i.e. STFT consistency, enforcement of the direct path, and speech magnitude constraint, are all instrumental in guiding BUDDy toward a solution with higher fidelity to clean speech, as measured by PESQ. We show DNS-MOS figures out of completeness. However, DNS-MOS variations are small across ablations and not indicative of fidelity to reference speech as DNS-MOS is not intrusive.
@@
Additionally, we examine the effect of parameterizing the likelihood model with a $L^{2}$-distance on compressed spectrograms rather than on waveforms as in previous work [26]. To do so, we replace the cost function $C(\cdot,\cdot)$ from (9), which is based on compressed STFT representations, with a simpler waveform-domain $L^{2}$-distance, and we empirically pick the optimal corresponding scaling factor $\tilde{\zeta}$. The results clearly show the superiority of the proposed cost function (9) using compressed STFT representations.
IV-A7 Listening experiment
@@
Instrumental metrics offer only limited insights into the performance of dereverberation algorithms [80]. We therefore conduct a listening experiment based on the MUSHRA recommendation [81] to assess the performance of BUDDy as perceived by human listeners. The test comprised 12 pages, featuring 6 reverberant speech utterances from the VCTK-RealReverb (matched) and the VCTK-SimulatedReverb (mismatched) sets. Participants were asked to rate the different stimuli with a single number representing overall quality, taking into account factors such as voice distortion, residual reverberation, and potential artifacts [80]. The test stimuli include our proposed method BUDDy, the unsupervised WPE [7] and RVAE-EM [38], as well as the supervised baselines PSE and SGMSE [24], Further details on the organization of the listening experiment are reported in Appendix -C1.
TABLE III: Ablation study on VCTK-RealReverb.
@@
  - RMS Power Constraint 2.22 $\pm$ 0.50 3.64 $\pm$ 0.50
    - Fixed Direct Path 2.10 $\pm$ 0.46 3.78 $\pm$ 0.44
      - STFT Consistency 1.96 $\pm$ 0.41 3.84 $\pm$ 0.39
@@
The results of the experiment are presented in Fig. 2. It can be observed that the unsupervised baselines WPE and RVAE-EM received low scores. Yet, RVAE-EM performs consistently better than WPE in this listening experiment, as opposed to what is suggested by instrumental metrics in Table I. In the matched test set (Fig. 2a), BUDDy obtained significantly lower scores than PSE and SGMSE+ ($p<0.001$ in a paired Welch test). However, in the mismatched set, PSE and SGMSE+ suffered a decrease in performance, losing up to $20$ points (out of $100$), while BUDDy mantained similar scores. In that case, there is no significant difference in performance between the three approaches ($p>0.1$), which closes the gap between BUDDy and the top-performing supervised baselines in this mismatched setting, highlighting the advantage provided by unsupervised learning.
Figure 2: Listening test results on VCTK-RealReverb and VCTK-SimulatedReverb. The boxplot shows first quartile, median, and third quartile.
IV-A8 Computational complexity
A current limitation of the proposed method lies in its high computational budget. Compared to classical diffusion approaches, such as SGMSE+ [24], our method requires several optimization iterations for the reverberation operator parameters at each reverse diffusion step, which increases the per-step inference time from 0.03s to 0.10s between SGMSE+ and BUDDy for each second of speech signal processed, as measured on a NVIDIA RTX A6000 GPU. Together with the larger number of reverse diffusion steps, the overall inference time is significantly higher than e.g. SGMSE+ or predictive approaches like PSE in Table I. Thus, while BUDDy is able to produce unprecedented high quality speech dereverberation without training on paired data, it cannot currently do so without an increase in computational complexity.
IV-B Singing Voice Dereverberation
@@
TABLE IV: Singing voice dereverberation results on NHSS dataset. Boldface numbers indicate best performance for supervised and unsupervised methods separately
Matched Mismatched
Method Unsup. $\ell^{1}$ STFT FAD $\ell^{1}$ STFT FAD
Reverberant - 1.98 $\pm$ 0.66 6.41 1.86 $\pm$ 0.56 5.65
PSE 1.56 $\pm$ 0.70 1.15 2.24 $\pm$ 0.78 1.79
SGMSE+ [24] 1.32 $\pm$ 0.51 0.82 1.37 $\pm$ 0.39 0.65
WPE [7] 2.02 $\pm$ 0.65 4.66 2.29 $\pm$ 0.66 5.74
Saito et al. [34] 1.95 $\pm$ 0.65 5.46 1.77 $\pm$ 0.52 4.90
BUDDy 1.90 $\pm$ 0.59 0.88 1.91 $\pm$ 0.50 0.60
@@
Figure 3: Listening test results on singing voice datasets NHSS-RealReverb and NHSS-SimulatedReverb. The boxplot shows first quartile, median, and third quartile.
IV-B1 Data
@@
We collect several publicly available singing voice datasets [82, 83, 84, 85, 86, 87]. These datasets feature over 94 h of studio-quality solo singing from a diverse array of singers and singing styles, spanning various languages. The majority of the recordings are in Chinese, followed by English, Japanese, and Korean. All datasets are down-sampled to 44.1 kHz. For testing, similar to [35], we use the sung part of NHSS [88, 89]. The NHSS dataset contains 100 English-language pop songs, 10 for each of the five male and five female singers recruited. We select a subset (90%) of the RIRs curated for the VCTK-based experiments, such that we only retain the RIRs whose original sample rate is at least 44.1 kHz. The resulting reverberant test set is referred to as NHSS-RealReverb. As for the speech voice experiments, we also prepare mismatched test set denoted as NHSS-SimulatedReverb using simulated RIRs sampled at 44.1 kHz.
IV-B2 Baselines
We evaluate the performance of BUDDy against two unsupervised baselines: WPE [7] and the unsupervised method from Saito et al. [34] which was originally designed for singing voice dereverberation. Additionally, we train supervised baselines following the same approach as in the speech dereverberation experiments described in Section IV-A2. Specifically, we compare BUDDy to PSE and SGMSE+, which share the same architectural design as BUDDy’s diffusion model, as detailed in Section IV-B3.
IV-B3 Hyperparameters
Similar to the speech dereverberation experiments, we adopt NCSN++ as the score model architecture. However, we adjust its hyperparameters to accommodate the higher sampling rate of 44.1 kHz. More details concerning the architecture and specific training configuration and inference hyperparameters are reported in Appendix -C2
IV-B4 Evaluation metrics